# Exploring Content and Social Connections of Fake News with Explainable Text and Graph Learning

### v1

@@ 3.1 Problem Statement
We define a post-node $\mathcal{P}_{i}=(\mathcal{T},\mathcal{G}_{k},\mathcal{M},lang)$, where $\mathcal{T}$ is the textual content; $\mathcal{G}_{k}$ defines the local $k$-hop network connections of the post, such as replies, quotes, and re-posts; $\mathcal{M}$ is the post’s shallow metadata set, which includes, for example, the tweet’s owner, the number of likes, the location, etc; and $lang$ is the language in which the tweet was originally written. Our goal is for a model $\sigma$ to identify and explain the most representative features within $\mathcal{P}_{i}$ that lead a post to be classified as misinformation (i.e., $y_{i}=\sigma(\mathcal{P}_{i})=0$) or fact (i.e., $y_{i}=\sigma(\mathcal{P}_{i})=1$).
@@
Post Encoding Module. This module comprises the encoding of a given post-node $\mathcal{P}_{i}$ in its vector representation. The module is divided into two components to represent two modalities: the post’s shallow metadata set $\mathcal{M}_{i}$; and its textual content $\mathcal{T}_{i}$. The first part obtains the vector representation of the shallow metadata features with the number of replies, quotes, and re-posts (Fig. 1). a) The second part encompasses a language proxy defined by $lang$ followed by the language-specific text encoders (Fig. 1. b) The language proxy acts as a router, assigning the post’s textual content $\mathcal{T}$ to a pretrained language-specific model (LM). The pre-trained language-specific model encoder generates a vector representation of the textual content. Finally, as described in Equation 1, both feature vectors are concatenated to form a unique multimodal post-node vector representation $\mathcal{X}_{\mathcal{P}_{i}}$.
@@ 3.3 Framework Instantiation
=${}_{\mathcal{P}_{i}}=\sigma\left(\textrm{{W}}\sum_{j\in\mathcal{G}_{k}}\alpha_{kj}\mathcal{X}_{\mathcal{P}_{j}}\right).$
@@ 4 Experiments
We carry out our evaluations on a dataset including multilingual and multi-topical texts and network data that contextualise the texts in the social media environment. As far as we know, the MuMiN [20] dataset is the only dataset that fulfils all requirements of our evaluation. In addition, our focus is on characterising the representation of the textual content and relational features. The MuMiN dataset is a publicly available dataset of multimodal data from Twitter. MuMiN links tweets covering multiple topics and languages with fact-checked claims, including text, metadata, and visual content from the tweets. For our study, we relied on the MuMiN-small version of this dataset, which consists of 2,1832,1832 , 183 claims and 7,202,5067,202,5067 , 202 , 506 tweets. To streamline our analysis, we filtered the dataset to consider tweets written in the three most prevalent languages (English, Portuguese, and Spanish), four types of entities (Claim, Tweet, Reply, and User), and six relations (Posted, Mentions, Retweeted, Quote_Of, Reply_To, and Discusses).
We developed our framework using PyTorch [22]. The full code of the framework, the described instantiation and example notebooks are available for review. The dataset split used was the same as proposed in the MuMiN dataset. As described in Section 3.3, our misinformation classifier is a GAT to account for the relational information. The classifier was trained using a single Nvidia RTX4070 Mobile GPU, with a learning rate of 0.0050.0050.005, 161616-dimensional hidden layer, and Adam optimiser [13] for a total of 400400400 epochs – the optimal setting we found in our experiments. The textual features were encoded using pre-trained language-specific models. The textual encoding is then submitted to a linear mapping of its original 768768768-dimensional space vector to a 812812812-dimensional space vector to match the dimensions of the shallow metadata set representation. The experiments in Sections 4.2-4.4 regard only the multilingual configuration, whilst comparing the various modalities.
@@ 4.1 Explainable Tweet Misinformation Detection
[…]
@@ 4.3 Trustworthiness of Explanations
[…]
Top-111 Top-222 Top-333 Top-555 Top-101010
@@
[…]

### v3

@@ 3.1 Problem Statement
We define a post-node $\mathcal{P}_{i}=(\mathcal{T},\mathcal{G}_{k},\mathcal{M},lang)$, where $\mathcal{T}$ is the textual content; $\mathcal{G}_{k}$ defines the local $k$-hop network connections of the post, such as replies, quotes, and re-posts; $\mathcal{M}$ is the post’s shallow metadata set, which includes, for example, the tweet’s owner, the number of likes, the location, etc; and langlang is the language in which the tweet was originally written. Our goal is for a model $\sigma$ to identify and explain the most representative features within $\mathcal{P}_{i}$ that lead a post to be classified as misinformation (i.e., $y_{i}=\sigma(\mathcal{P}_{i})=0$) or fact (i.e., $y_{i}=\sigma(\mathcal{P}_{i})=1$).
@@
Post Encoding Module. This module comprises the encoding of a given post-node $\mathcal{P}_{i}$ in its vector representation. The module is divided into two components to represent two modalities: the post’s shallow metadata set $\mathcal{M}_{i}$; and its textual content $\mathcal{T}_{i}$. The first part obtains the vector representation of the shallow metadata features with the number of replies, quotes, and re-posts (Fig. 1). a) The second part encompasses a language proxy defined by langlang followed by the language-specific text encoders (Fig. 1. b) The language proxy acts as a router, assigning the post’s textual content $\mathcal{T}$ to a pretrained language-specific model (LM). The pre-trained language-specific model encoder generates a vector representation of the textual content. Finally, as described in Equation 1, both feature vectors are concatenated to form a unique multimodal post-node vector representation $\mathcal{X}_{\mathcal{P}_{i}}$.
@@ 3.3 Framework Instantiation
$_{\mathcal{P}_{i}}=\sigma\left(\textrm{{W}}\sum_{j\in\mathcal{G}_{k}}\alpha_{kj}\mathcal{X}_{\mathcal{P}_{j}}\right).$
@@ 4 Experiments
We carry out our evaluations on a dataset including multilingual and multi-topical texts and network data that contextualise the texts in the social media environment. As far as we know, the MuMiN [20] dataset is the only dataset that fulfils all requirements of our evaluation. In addition, our focus is on characterising the representation of the textual content and relational features. The MuMiN dataset is a publicly available dataset of multimodal data from Twitter. MuMiN links tweets covering multiple topics and languages with fact-checked claims, including text, metadata, and visual content from the tweets. For our study, we relied on the MuMiN-small version of this dataset, which consists of $2,183$ claims and $7,202,506$ tweets. To streamline our analysis, we filtered the dataset to consider tweets written in the three most prevalent languages (English, Portuguese, and Spanish), four types of entities (Claim, Tweet, Reply, and User), and six relations (Posted, Mentions, Retweeted, Quote_Of, Reply_To, and Discusses).
We developed our framework using PyTorch [22]. The full code of the framework, the described instantiation and example notebooks are available for review. The dataset split used was the same as proposed in the MuMiN dataset. As described in Section 3.3, our misinformation classifier is a GAT to account for the relational information. The classifier was trained using a single Nvidia RTX4070 Mobile GPU, with a learning rate of $0.005$, $16$-dimensional hidden layer, and Adam optimiser [13] for a total of $400$ epochs – the optimal setting we found in our experiments. The textual features were encoded using pre-trained language-specific models. The textual encoding is then submitted to a linear mapping of its original $768$-dimensional space vector to a $812$-dimensional space vector to match the dimensions of the shallow metadata set representation. The experiments in Sections 4.2-4.4 regard only the multilingual configuration, whilst comparing the various modalities.
@@ 4.1 Explainable Tweet Misinformation Detection
The empirical results of tweet classification are reported in Table 1, in which the model assigns each input information to the misinformation or fact label. We assess three different inputs: graph-based features, text-based features, and multimodal features, which are the concatenation of both graph-based and text-based features, and measure the obtained F1-score. The reported results are generated by $1000$ bootstrap resampling with $95\%$ confidence interval.
@@ 4.3 Trustworthiness of Explanations
Modality F1-Score given top-$k$ most relevant features
Top-11 Top-22 Top-33 Top-55 Top-$10$
@@
Figure 4: F1-Score of trustworthiness box-plot visualisation. Lines are the average F1-Score in each Top-$K$.
The protocol consists of, first, simulating the trust in individual predictions. For that, we randomly select $30\%$ of the features and assume these as “untrustworthy”, i.e., the user would likely recognise these features and would not want them as explanations. The second part of the protocol embraces the constructions of an oracle “trustworthiness”. We labelled the classifier’s prediction as untrustworthy if the prediction changes after removing the features selected as untrustworthy from an instance, and trustworthy otherwise. The oracle assumes that if removing untrustworthy features changes a prediction, those features are decisive for the classifier. In this simulated setup, the oracle serves as the true label for evaluating prediction credibility. The third step simulates user judgment by considering a prediction untrustworthy if it changes after removing all untrustworthy features identified in the explanation using a linear approximation model. Finally, we compare the oracle’s label with the simulated users’ decisions.
Given the trustworthiness evaluation protocol, we report and compare the F1-score on the trustworthy predictions for the classifier trained with graph-based features, text-based features, and multimodal features over 25 rounds, considering the top-$K$ most relevant explainable features ($K\in\{1,2,3,5,10\}$). The results are compiled in Fig. 4 and Table 3. The multimodal classifier achieved the most stable results, having the best F1-score in top-$1$ and top-$10$, whilst being the second-best in the other cases. The classifier that was solely trained with graph-based features displayed the worst results, indicating that it usually mistrusts trustworthy predictions, whilst trusting untrustworthy predictions too often. Despite the used modality, the results show F1-Score above $80\%$ (except the graph-based top-$1$), confirming the reliability of the mu2X explanations.